OpenStudy (anonymous):

find the solutions for 1/x^2 + 8/x +15

OpenStudy (gorv):

\[\frac{ 1+8x+15x^{2} }{ x^{2} }\]

OpenStudy (gorv):

\[\frac{ 15*x^{2} +5*x+3*x+1}{ x^{2} }=\frac{ (5x+1)*(3x+1) }{ x^{2} }\]

OpenStudy (anonymous):

ahhh - when I did it, I forgot to keep the x^2 in the denominator. thank you!!